OpenStudy (anonymous):

a line passes through the points (6, -2) and (5,-1) what is the x- intercept of the line ? who can show work for me please can you? i need help on this one @aum

OpenStudy (anonymous):

First, you know that the line has an equation : \[y=mx+b\] So you have to find m and b. for m we have : \[m=\frac{y_2-y_1}{x_2-x_1}\] Where : \[(x_1,y_1)\text{ and }(x_2,y_2)\] are two points from the line. Can you calculate m ?

OpenStudy (anonymous):

thanks i get that part now (: whats next ? how can i calculate m ?

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

The exercise says : "a line passes through the points (6, -2) and (5,-1)" So \[(x_1,y_1)=(6,-2)\text{ and } (x_2,y_2)=(5,-1).\] Did you get it ? Please, try to calculate m :)

OpenStudy (anonymous):

whats the answer then ?

OpenStudy (anonymous):

@Noura11

OpenStudy (shinalcantara):

First thing to consider is the slope of the line to get the equation \[m = \frac{ y_2-y_1 }{ x_2-x_1 }\] \[m = \frac{ -1-(-2) }{ 5-6 }\] \[m = -1\] ------------------------ Plug in to point slope form using one point. Let's have (6,-2) Point-slope form: \[(y-y_1) = m(x-x_1)\] \[y - (-2) = -1(x-6)\] \[y+2 = -x +6\] \[y = -x + 6 - 2\] \[y = -x +4\] ----------------------- Since you are looking for the x-intercept, note that y=0 y = -x + 4 0 = -x + 4 x = 4

OpenStudy (anonymous):

thankyouuuuuu guys for helping !!

OpenStudy (anonymous):

@shinalcantara @Noura11